#1. Consider the information about the economy of Pakistan. Note that the currency of Pakistan is the rupee.

The government purchases: 2.602.60 trillions of rupees.
Individuals consume: 8.508.50 trillions of rupees.
Individuals save: 5.355.35 trillions of rupees.
Businesses invest: 1.401.40 trillions of rupees.
Foreigners spend: 0.530.53 trillions of rupees to purchase Pakistani firms.
Pakistan imports: 2.492.49 trillions of rupees.
Pakistan exports: 1.211.21 trillions of rupees.

Calculate Pakistan's GDP. Assume that the values are all current and no conversions need to be made. Give your answer in terms of trillions of rupees, and round to two decimals.

Pakistans GDP: ________ trillions of rupees.

#2.

In the United States, illegal activities are excluded from calculations of GDP. That is not the case in every country: "Italy includes cocaine, prostitution and black market alcohol in GDP figures." (The Independent, October 30, 2014)

How would you expect the inclusion of these underground activities to impact the value of Italy's GDP?

The value of GDP will?

- rise, fall or will not change?

Which statements provide evidence that GDP imperfectly reflects the well‑being of a nation?

GDP calculations do not: (select as many as possible)

A. value production that is not intended for sale.

B. put importance on physical health.

C. include the cost of intermediate goods.

D. consider the quality of the environment.
Answer #1

solution:

1. GDP= C+ I+ G+ NX where c= consumption, I= investment, G= government expenditure and N X= net exports

thus Pakistan's GDP= 8.50+1.40+2.60+(1.21 -2.49)= 11.22 trillion rupees.

Individual saving and foreigners spending are redundant items. Hence not used in calculation of GDP.